A car is travelling with a constant speed of 80km/h and passes a stationary motorcycle policeman. The policeman sets off in pursuit, accelerating to 80km/h in 10  minutesreaching a constant speed of 100 km/h after a further 5 seconds. At what time will the policeman catch up with the car?



 

Grade:11

2 Answers

Vikas TU
10 years ago

a) Distance travelled by the car = 80,000/3600 x 10 x 60 = 1333.3 m

b) Distance travelled by police car in 10 minutes = 80,000/3600 x 5 = 111.1 m

    Distance travelled by police car in ''t'' seconds = 100,000/3600 x t = 27.77t m

c) Now Distance travelled by police car = Distance travelled by car
   [ IN  ORDER  TO  CATCH  THE CAR  DISTANCE SHOULD BE EQUAL]

 

d)  1333.3 = 111.1 + 27.77t

    27.77t  = 1222.2

t = 44.01 seconds

 

Thus 

the policeman will catch up the car after 44.01 seconds.

 

Plz Approve!

:)

 

V. Harshit -
8 years ago
thank you